# Few-Shot Domain Adaptation for Learned Image Compression

### v1

@@ IV-B Adaptation Performance
IV-B1 Comparisons with Traditional Codec
We compare the proposed method with the conventional standard VVC by VTM-22.0 intra coding. We plot the rate-distortion (RD) curves on Kodak (in-domain) and the other five domains in Fig. 6. All pre-trained models perform comparably or better than VTM on Kodak, but most of them fall behind VTM on the other five domains. Our method enhances these models to the level of VTM except on the Pixel domain, where the domain gap is too wide to fully compensate [32].
TABLE III: Detailed BD-rate ($\downarrow$) using different numbers of target samples ($N$) for domain adaptation. We compare our method with full-model finetune and DANICE [24]. The anchors are corresponding pre-trained models. Note that DANICE can not be applied to ELIC as GDN is not used.
@@ IV-B1 Comparisons with Traditional Codec
IV-B2 Comparisons with Domain Adaptation Method
To evaluate the efficiency, we also compare with existing DA method DANICE [24] and full-model finetune by RD performance and complexity. We demonstrate detailed adaptation results under different $N$ by computing BD-rate [7] in Table III. It is noted that DANICE is restricted to GDN structures [3], making it inapplicable to baselines like ELIC. Additionally, DANICE fails to achieve domain adaptation in some conditions, such as TCM and MLIC++ on the Game domain, indicating a limited versatility. Our method, however, achieves excellent adaptation performance in all conditions, performing comparably with full-model finetune and significantly outperforming DANICE.
In the regard of complexity, we highlight the compactness of the proposed adapters in Table I and Table II. Our method only needs to transmit fewer than $2\%$ of the total parameters, while much more parameters need considering in DANICE and finetune due to the complex entropy model. Besides, we examine the variations in decoding time and FLOPs introduced by the proposed adapters in Table II. Our method increase FLOPs by less than 10% and coding time by less than 1.5%.
Figure 7: Compatibility between our method and representative instance adaptation method [32] (abbreviated as DA and IA, respectively). We plot the rate-distortion curves on the Pixel and Screen-content domain.
IV-B3 Comparisons with Instance Adaptation Method
We compare our proposed domain adaptation (DA) method for LIC with existing instance adaptation (IA) methods [10; 43; 42; 38; 32]. IA adapts a pre-trained LIC model to individual images through per-image online training during inference. In contrast, DA involves a one-time training process on multiple images, adapting the model to a target domain before inference. Consequently, we suggest that our proposed DA method for LIC complements existing IA methods. The compatibility between these methods is illustrated in Fig. 7.
Specifically, we explore the combination of our proposed method with one of the latest IA methods [32]. We apply both techniques to the Pixel and Screen-content domain. For alignment, we use the pre-trained Cheng2020 model [11] as the baseline. Our DA method is conducted with 25 target samples, followed by implementing IA for each image on the adapted model. The rate-distortion curves, compared with VTM-22.0 intra coding, the pre-trained Cheng2020, and each technique individually, are shown in Fig. 7. Note that some curves differ slightly due to the different padding method adopted by Lv et al. [32] in their released code; we respect their implementation in Fig. 7. When combining both techniques, the RD performance is further improved, indicating that our method complements IA by introducing additional domain information from a limited number of samples.
Regarding encoding complexity, IA relies on expensive per-image online training, whereas DA is more efficient due to its one-time training on multiple samples. We display the encoding time for processing different numbers of test images in Fig. 8. For DA, a fixed one-time training is required to adapt the pre-trained LIC model to the target domain. Adapting the pre-trained Cheng2020 model with 25 samples takes less than half an hour on a single Tesla V100 GPU. Once adapted, the model achieves similar encoding speed on test images as the source model. In contrast, IA requires time-consuming online training for every test image, resulting in a significantly longer per-image encoding time compared to the source model. Therefore, DA is considerably more efficient than IA when encoding a batch of target-domain images.
IV-C Ablation Studies
TABLE IV: Ablation studies conducted on ELIC adapted with 25 samples. The anchor of BD-rate ($\downarrow$) is pre-trained ELIC models.
Stage 1 Stage 2 Pixel Screen-content Craters Game Pathology AVG
Conv LoRA Stack 1234 Stack 4
-19.38% -23.73% -10.84% -12.73% -10.30% -15.40%
-20.40% -24.28% -12.57% -13.40% -11.50% -16.43%
-19.92% -22.46% -11.95% -10.68% -10.95% -15.19%
-23.35% -24.95% -13.77% -13.83% -12.29% -17.64%
TABLE V: Explorations on the structures of Conv-Adapters. Evaluated by the parameters to transmit and the BD-rate ($\downarrow$).
@@ IV-C Ablation Studies
TABLE VI: Explorations on the deployment of Conv-Adapters. Evaluated by the parameters to transmit and the BD-rate ($\downarrow$).
@@
Figure 8: Encoding time of different adapation methods when applied to different numbers of test images from the Pixel domain, using a single Tesla V100 GPU. The proposed few-shot domain adaptation method is conducted on 25 target samples.
We conduct ablation studies to evaluate the contributions of Conv-Adapters, LoRA-Adapters and the two-stage training strategy. Using pre-trained ELIC as the baseline, we perform the proposed adaptation with 25 samples on the five domains. The results are shown in Table IV. Conv-Adapters achieve an average reduction of 15.40% in BD-rate compared to pre-trained models, while LoRA-Adapters and the two-stage training strategy each contributes an additional 1% BD-rate reduction. We also compare our second stage strategy with finetuning all Conv-Adapters in the decoder and display the superiority of our method.
Given that Conv-Adapter, which involves linear re-allocation on channels, has the most significant impact on adaptation performance, we further explore the structures of Conv-Adapters in Table V and the deployment strategy in Table VI. We conduct adaptation on ELIC using 25 target samples and evaluate the performance based on the parameters to transmit, and the BD-rate between the adapted model and the source model. Our results demonstrate the efficiency of our method in terms of both BD-rate and parameter requirements.
For the structures, we first replace the proposed Conv1$\times$1 with more compact structures such as GDN [3] and depthwise Conv3$\times$3. We then test more complex structures, including Conv3$\times$3 and a sequence of depthwise Conv3$\times$3 followed by Conv1$\times$1. For all structures, we follow the restriction on initialization and initialize them to perform identical transforms at the start of training. Using ELIC [19] as the baseline model, we conduct adaptation experiments on five domains, each with 25 target samples. The training strategy for adapters remains consistent. As shown in Table V, compact structures like GDN and depthwise Conv3$\times$3 have fewer parameters to transmit but exhibit significant degradation in average adaptation performance. For more complex structures, the increased parameters in the spatial dimension fail to further improve performance, underscoring the efficiency of Conv1$\times$1 in balancing complexity and performance.
For the deployment strategy, following Fig. 5, we divide the analysis transform and the synthesis transform into four stacks and gradually apply Conv-Adapters to each stack. Although the parameters to transmit increase, applying Conv-Adapters to all stacks significantly enhances adaptation performance. Therefore, we recommend deploying Conv-Adapters in all stacks. We visualize the kernels after training in Fig. 10.
@@
IV-D Analyses on Channel-wise Re-allocation
IV-D1 Variation on Channels
We use the output latent of the analysis transform, and study the distribution of logarithmic energy and entropy on channels following [19] in Fig. 11. The results are averaged on the Pixel domain. We find that both energy and entropy on specific channels drop after adapting with adapters. Especially, the variation is more apparent on low-energy channels, while high-energy channels remain almost unchanged. This suggests that the energy becomes more concentrated to high-energy channels on out-of-domain images.
IV-D2 Reconstructions from Low-Energy Channels
To further investigate the variation on channel-wise decomposition, we compare the reconstructions from low-energy channels before and after adaptation. Concretely, we keep specific channels unchanged and mask the rest into zeros, reconstruct the modified latent, then subtract the reconstruction of all-zero one. We sort the channels by energy, and display the reconstructions from specific ranges of channels after $200th$ in Fig. 9, which is similar to Fig. 3. We find weaker details of inputs can be observed in reconstructions from the model adapted with our method, suggesting high-frequency components have been better moved out of the low-energy channels, which is more in line with the compact energy allocation in processing in-domain images. Thus, we suggest the pre-trained frequency decomposition on high-frequency components is strengthened by our method.
Figure 12: Reconstructions and PSNR ($\uparrow$) /bpp ($\downarrow$) performance compared with ground truth and the pre-trained model, detailed by specific $50\times 50$ crops. The reconstruction errors with ground truth are shown in right.
IV-E Visualization of Reconstructions
We present the qualitative reconstruction results of the pre-trained ELIC [19], the model adapted with our method using 25 samples, and the corresponding reconstruction errors in Fig. 12. The model is optimized using MSE with $\lambda=0.483$. We showcase out-of-domain samples from the five domains we used: Pixel, Screen-content, Craters, Games, and Pathology. Specific $50\times 50$ crops from the reconstructions are displayed with pixel enlargement for better visibility.
In addition to RD improvement, reconstructions from the model adapted with our method demonstrate superior visual quality, particularly in high-frequency edges and details. For Pixel and Screen-content samples, the sharp edges of pixel-style art and letters are reconstructed more accurately. For samples from Craters, Games, and Pathology, our reconstructions capture more details of the ground truth. Additionally, we compute the reconstruction errors relative to the ground truth, finding that our method primarily reduces reconstruction errors at edges and in fine details.
V Conclusion

### v2

@@ Appendix B Detailed Domain Adaptation Performance
Table 4: Complexity compared with source models. The results are averaged on Kodak with a single Tesla V100 GPU.
Baseline Finetune Ours
Cheng2020 0.246 0.216
ELIC 0.418 0.356
TCM 0.672 0.536
MLIC++ 0.574 0.512
Table 5: Comparisons of training speed (second per iteration) between our method and full-model finetune.
@@
Table 6: Detailed BD-rate ($\downarrow$) using different numbers of target samples ($N$) for domain adaptation.
Appendix C Computational Complexity
In this section, we examine the variations in decoding time and FLOPs introduced by the proposed adapters. The results are presented in Table 4. Compared to the source models, our adapters increase FLOPs by less than 10% and coding time by less than 1.5%. Additionally, we provide the training speed for the four baselines we used, mainly comparing our method with full-model fine-tuning. Using $256\times 256$ patches as input and a batch size of 4, the results are shown in Table 5, obtained with a single Tesla V100 GPU. The training speeds of both methods are of the same magnitude, with our method being slightly faster.
Appendix D Explorations on Structure and Deployment of Conv-Adapters
We further explore the structure of Conv-Adapters in Table 7 and the deployment strategy in Table 8. We conduct adaptation on ELIC using 25 target samples and evaluate the performance by parameters and rate-distortion improvement.
@@ Appendix D Explorations on Structure and Deployment of Conv-Adapters
Table 7: Explorations on the structures of Conv-Adapters. Evaluated by the parameters to transmit and the BD-rate ($\downarrow$).
@@
Table 8: Explorations on the deployment of Conv-Adapters. Evaluated by the parameters to transmit and the BD-rate ($\downarrow$).
For structure, we first replace the proposed Conv1$\times$1 with more compact structures such as GDN (Ballé, Laparra, and Simoncelli 2015) and depthwise Conv3$\times$3. We then test more complex structures, including Conv3$\times$3 and a sequence of depthwise Conv3$\times$3 followed by Conv1$\times$1. For all structures, we follow the restriction on initialization and initialize them to perform identical transforms at the start of training. Using ELIC (He et al. 2022) as the baseline model, we conduct adaptation experiments on five domains, each with 25 target samples. The training strategy for adapters remains consistent. As shown in Table 7, compact structures like GDN and depthwise Conv3$\times$3 have fewer parameters to transmit but exhibit significant degradation in average adaptation performance. For more complex structures, the increased parameters in the spatial dimension fail to further improve performance, underscoring the efficiency of Conv1$\times$1 in balancing complexity and performance.
For deployment, following Fig. 5, we divide the analysis transform and the synthesis transform into four stacks and gradually apply Conv-Adapters to each stack. Although the parameters to transmit increase, applying Conv-Adapters to all stacks significantly enhances adaptation performance. Therefore, we recommend deploying Conv-Adapters in all stacks. We visualize the kernels after training in Fig. 11.
Appendix E Visualization of Reconstructions
We present the reconstructions of the pre-trained ELIC (He et al. 2022), the adapted ELIC using 25 samples, and the corresponding reconstruction errors in Fig. 13. The model is optimized using MSE with $\lambda=0.483$. We showcase out-of-domain samples from the five domains we used: Pixel, Screen-content, Craters, Games, and Pathology. Specific $50\times 50$ crops from the reconstructions are displayed with pixel enlargement for better visibility. In addition to RD improvement, reconstructions from the adapted model demonstrate superior visual quality, particularly in high-frequency edges and details. For Pixel and Screen-content samples, the sharp edges of pixel-style art and letters are reconstructed more accurately. For samples from Craters, Games, and Pathology, our reconstructions capture more details of the ground truth. We also compute the reconstruction errors relative to the ground truth, suggesting our method primarily reduces reconstruction errors at edges and in fine details.
@@ Appendix E Visualization of Reconstructions
Figure 12: Performance compared with official results.
Appendix F Details for Reproducing ELIC
We utilize DIV2K (Agustsson and Timofte 2017) and CLIC 2020 Professional (CLIC 2020) for training ELIC (He et al. 2022) models. We construct our code on the base of (Jiang 2022). During training, images are randomly cropped into 256 × 256 patches, and a batch size of 8 is adopted following official implementation. We optimize the model using distortion terms of MSE and set the Lagrange multiplier $\lambda\in\left\{18,35,67,130,250,483\right\}\times 10^{-4}$ for different quality presets. We use Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$. We train each model for 2M steps. The learning rate starts from $10^{-4}$ and drops to $3\times 10^{-5}$ at 1.2M steps, then to $10^{-5}$ at 1.5M steps, and finally to $10^{-6}$ at 1.8M steps. The performance of our models is generally consistent with the official-released results as shown in Fig. 12. It is noted that a different set of $\lambda$ was adopted in official results, where the 6 median values were $\left\{16,32,75,150,300,450\right\}\times 10^{-4}$.
Figure 13: Reconstructions and PSNR ($\uparrow$) / bpp ($\downarrow$) comparisons, detailed by specific 50 × 50 crops. The reconstruction errors with ground truth are shown in right.